# Data-driven decision-making under uncertainty with entropic risk measure

### v1

@@ 4.1 Entropic Risk Matching
In this section, our goal is to learn a GMM whose samples mimic the bias in the loss scenarios. This bias corresponding to the learned GMM is estimated using bootstrapping and added to the empirical entropic risk to mitigate the underestimation problem.
The objective of Algorithm 2 is to learn the parameters of a GMM, denoted $\mathbb{Q}^{{{\boldsymbol{\theta}}}}$, which minimize the Wasserstein distance between two distributions. The first distribution is one of the entropic risk derived from $N$ i.i.d. samples from $\mathbb{Q}^{{{\boldsymbol{\theta}}}}$, while the second distribution is the distribution of the entropic risk computed using a subset of the observed loss scenarios from the true distribution $\mathbb{P}$.
Initially, the algorithm starts by partitioning the given loss scenarios, denoted as $\mathcal{X}=\{\ell(\boldsymbol{\xi}_{1}),\ell(\boldsymbol{\xi}_{2}),\dots,\ell(\boldsymbol{\xi}_{N})\}$ into $B$ bins, where $\ell(\boldsymbol{\xi}_{i})$ represents the loss corresponding to the random variable $\boldsymbol{\xi}_{i}$. Each bin contains $n=N/B$ loss scenarios. For each bin, the entropic risk is computed, producing an empirical distribution ^$\hat{\mathbb{P}}_{n}^{\rho}$ of the empirical entropic risk based on the observed data. This empirical distribution represents the entropic risk derived directly from the true loss scenarios.
Next, we fit a GMM, referred to as ^$\hat{\mathbb{Q}}^{{\boldsymbol{\theta}}}$ to the loss scenarios in $\mathcal{X}$. The GMM is fitted using the Expectation-Maximization algorithm, which provides initial estimates of the model parameters $\boldsymbol{\theta}=(\hat{\boldsymbol{\pi}},\hat{\boldsymbol{\mu}},\hat{\boldsymbol{\sigma}})$. These parameters include the mixing weights $\hat{\boldsymbol{\pi}}$, means $\hat{\boldsymbol{\mu}}$, and standard deviations $\hat{\boldsymbol{\sigma}}$ for each component of the GMM. To compare the model-based distribution of entropic risk to the empirical distribution ^$\hat{\mathbb{P}}_{n}^{\rho}$ of entropic risk, samples are generated from the GMM. Specifically, $B^{\prime}\times n$ i.i.d. samples are drawn from the initial GMM with parameters $\boldsymbol{\theta}$, and these samples are divided into $B^{\prime}$ bins. The entropic risk is computed for each bin, producing a model-based distribution ^$\hat{\mathbb{Q}}_{n}^{\rho,{{\boldsymbol{\theta}}}}$ of the entropic risk. This distribution is derived from the GMM and is used to assess how well the GMM replicates the risk profile of the true data.
@@
The next step is to compare the empirical distribution ^$\hat{\mathbb{P}}_{n}^{\rho}$ with the model-based distribution ^$\hat{\mathbb{Q}}_{n}^{\rho}$ using the Wasserstein distance, denoted as $\mathcal{W}^{p}\left(\hat{{\mathbb{P}}}_{n}^{\rho},\hat{\mathbb{Q}}_{n}^{\rho,{{\boldsymbol{\theta}}}}\right)$, where the Wasserstein distance is defined as follows:
$\mathcal{W}^{p}\left({\mathbb{P}}_{1},{\mathbb{P}}_{2}\right)=\mathop{\rm inf}_{\pi\in\mathcal{M}(\Xi\times\Xi)}\left\{\left(\int_{\Xi}\int_{\Xi}\|{{{\boldsymbol{\zeta}}}}_{1}-{{{\boldsymbol{\zeta}}}}_{2}\|^{p}\,\pi(d{{\boldsymbol{\zeta}}}_{1},d{{\boldsymbol{\zeta}}}_{2})\right)^{1/p}\right\}$ (6)
@@
where $\pi$ is a joint distribution of ${{{\boldsymbol{\zeta}}}}_{1}$ and ${{{\boldsymbol{\zeta}}}}_{2}$ with marginals ${\mathbb{P}}_{1}$ and ${\mathbb{P}}_{2}$, respectively. This distance quantifies the discrepancy between the two distributions. The algorithm iteratively adjusts the GMM parameters to reduce this distance, using gradient descent to update the parameters $\boldsymbol{\theta}_{t}$ at each iteration. Specifically, the parameters are updated according to:
@@
$\boldsymbol{\theta}_{t+1}=\boldsymbol{\theta}_{t}-\gamma\nabla_{\boldsymbol{\theta}_{t}}\mathcal{W}^{p}\left(\hat{{\mathbb{P}}}_{n}^{\rho},\hat{\mathbb{Q}}_{n}^{\rho,{{\boldsymbol{\theta}}}_{t}}\right)$
where $\gamma$ is the step size. Algorithm 6 generates “differentiable samples” from the GMM, $\mathbb{Q}^{{{\boldsymbol{\theta}}}_{t}}$, which are then used to compute the Wasserstein distance. This enables the calculation of the gradients of the Wasserstein distance with respect to ${{\boldsymbol{\theta}}}_{t}$ (see Appendix B).
After each update, the parameters are projected back into the feasible region for a valid GMM. This ensures that the mixing weights $\hat{\boldsymbol{\pi}}_{t+1}$ remain valid probabilities (which is achieved using a softmax function), and that the standard deviations $\hat{\boldsymbol{\sigma}}_{t+1}$ are positive (enforced by taking the maximum of $\exp(-5)$ and $\hat{\boldsymbol{\sigma}}_{t+1}$).
The iterative process continues until the Wasserstein distance $\mathcal{W}^{p}\left(\hat{{\mathbb{P}}}_{n}^{\rho},\hat{\mathbb{Q}}_{n}^{\rho,{{\boldsymbol{\theta}}}_{t}}\right)$ falls below a predefined convergence threshold $\epsilon$, or until a maximum number of iterations $T$ is reached. Once convergence is achieved, the final GMM, $\mathbb{Q}_{N}$, is returned as the learned model that best approximates the entropic risk distribution of the true data. The number of components $Y$ in the GMM is selected by choosing the value that minimizes the Wasserstein distance. By iteratively adjusting the parameters and refining the number of components, the algorithm ensures that the learned GMM closely matches the distribution of entropic risk from the original data.
@@
1:function BS-Match($\mathcal{X},Y$)
2:     Divide loss scenarios in $\mathcal{X}$ into $B$ bins, each of size $n$
3:     Compute the entropic risk in $B$ bins, forming the empirical distribution ^$\hat{{\mathbb{P}}}_{n}^{\rho}$
4:     Fit an GMM $\mathbb{Q}^{{\boldsymbol{\theta}}}$ to $\mathcal{X}$ with Y components using EM algorithm
5:     Initialize the iteration counter $t\leftarrow 0$ and $d\leftarrow\infty$
@@
8:         Compute entropic risk in each bin, forming ^$\hat{\mathbb{Q}}_{n}^{\rho,{{\boldsymbol{\theta}}}_{t}}$
9:         Update GMM parameters: ${{\boldsymbol{\theta}}}_{t+1}\leftarrow{{\boldsymbol{\theta}}}_{t}-\gamma\nabla_{{{\boldsymbol{\theta}}}_{t}}\mathcal{W}^{p}\left(\hat{{\mathbb{P}}}_{n}^{\rho},\hat{\mathbb{Q}}_{n}^{\rho,{{\boldsymbol{\theta}}}_{t}}\right)$
@@
11:         Update distance: $d\leftarrow\mathcal{W}^{p}\left(\hat{{\mathbb{P}}}_{n}^{\rho},\hat{\mathbb{Q}}_{n}^{\rho,{{\boldsymbol{\theta}}}_{t}}\right)$
@@
14:     $\mathbb{Q}_{N}\leftarrow\mathbb{Q}^{{{\boldsymbol{\theta}}}_{t}}$
15:     return $\mathbb{Q}_{N}$
16:end function
Once we learn the distribution $\mathbb{Q}_{N}$ using Algorithm 2, we use it in Algorithm 1 to estimate the bias
$\delta_{N}(\mathbb{Q}_{N}):=\mbox{median}[\rho_{\mathbb{Q}_{N}}(\zeta)-\rho_{\hat{\mathbb{Q}}_{N,N}}(\zeta)|\mathbb{Q}_{N}].$
Even though computing the Wasserstein distance between distribution of losses has a worst-case complexity $O(B^{\prime}\log(B^{\prime}))$ (Kolouri et al.,, 2018), there is a significant total cost associated with the gradient descent procedure described in Algorithm 2. In the next section, we provide a semi-analytic procedure to learn a two-component GMM that can account for the tail scenarios.
@@ 4.2 Matching the extremes
This section aims to approximate the tails of the true underlying distribution in order to improve the estimation of the entropic risk, which depends on the full distribution but is particularly sensitive to the tails. Our approach uses the distribution of tails of normal distribution to approximate the distribution of the maxima of the data, motivated by both extreme value theory and the need for an analytical solution.
@@
Consider a random variable $\zeta\sim F(\cdot)$. Given i.i.d. samples of $\{\zeta_{1},\zeta_{2},\cdots,\zeta_{n}\}$ drawn from $F(\cdot)$, the maximum of the samples is denoted as $M_{n}=\mathop{\rm max}\{\zeta_{1},\zeta_{2},\cdots,\zeta_{n}\}$. According to the Fisher–Tippett–Gnedenko theorem, the distribution of a (normalized) maxima $M_{n}$ converges to a non-degenerate distribution $G$:
$\lim_{n\rightarrow\infty}{\mathbb{P}}\left(\frac{M_{n}-b_{n}}{a_{n}}\leq x\right)=\lim_{n\rightarrow\infty}F(a_{n}x+b_{n})^{n}\rightarrow G(x),$ (7)
@@
where $a_{n}$ and $b_{n}$ are normalizing sequences of scale and location parameters, respectively, that ensure the limit exists. The limit distribution $G$ belongs to one of three extreme value distributions–Weibull, Fréchet or Gumbel distribution–depending on the tails of $F(\cdot)$ (Haan and Ferreira,, 2006).
While the entropic risk depends on the entire distribution of losses, its sensitivity to extreme values makes it important to approximate the tail behavior accurately. However, rather than directly fitting an extreme value distribution to the empirical tails, we approximate the distribution of the maxima by fitting a normal distribution and use this to construct a GMM. This approach is motivated by the fact that GMM provides a closed-form analytical solution for the entropic risk, making computations efficient in the bootsrapping step.
The key idea is not to fit a normal distribution to the data itself but rather to match the distribution of the maxima of $n$ i.i.d samples drawn from a normal distribution with the distribution of maxima of $n$ i.i.d samples drawn from the empirical data. This approach is inspired by the Fisher–Tippett–Gnedenko theorem, which indicates that the distribution of the normalized maxima will converge to one of the three possible extreme value distributions. By using a normal distribution, we strike a balance between capturing the tail behavior and retaining analytical tractability.
To empirically estimate the distribution of maxima from the observed data, we use the block maxima method, commonly applied in extreme value analysis (Haan and Ferreira,, 2006). The algorithm for fitting a normal distribution whose tails match those of the loss scenarios is outlined in Algorithm 3. First, we divide the scenarios in the set $\mathcal{X}$ into $B$ bins, each of size $n=N/B$ and then compute the maximum within each bin. Let $F_{n}(.)$ denote the cumulative distribution function (cdf) of the maxima.
The distribution of the maxima of $n$ i.i.d samples from a normal distribution $\mathcal{N}(\mu,\sigma)$ is given by $\left(\Phi_{\mu,\sigma}\right)^{n}$ where $\Phi_{\mu,\sigma}$ is the cdf of a normally distributed random variable with mean $\mu$ and $\sigma$. We find the parameters of the normal distribution by matching the 50th and 90th quantiles of $F_{n}(.)$ to the corresponding quantiles of $\left(\Phi_{\mu,\sigma}\right)^{n}$:
$\mu+\sigma\Phi_{0,1}^{-1}(0.5^{1/n})=F_{n}(0.5)$
$\mu+\sigma\Phi_{0,1}^{-1}(0.9^{1/n})=F_{n}(0.9),$
For analytical tractability in risk estimation, we restrict the model to a two-component GMM with equal weights for each component. The GMM’s mean is set to be the mean of the data $\hat{\mu}_{\mathcal{X}}$, and the variance of one component is set to zero, resulting in a GMM with one component as a Dirac delta distribution $\mathcal{N}(\mu_{0},0)$ and the other as a normal distribution $\mathcal{N}(\mu,\sigma)$. These components satisfy the condition: $0.5\mu_{0}+0.5\mu=\hat{\mu}_{\mathcal{X}}$.
@@
1:function BS-EVT($\mathcal{X}$)
2:     Divide scenarios in $\mathcal{X}$ into $B$ bins of size $n=N/B$ each
3:     Store maximum of each bin in set $\mathcal{M}$
4:     $F_{n}\leftarrow$ cdf of maxima in $\mathcal{M}$
5:     Find $(\mu,\sigma)$ by matching $50^{\text{th}}$ and $90^{\text{th}}$ quantiles of $\left(\Phi_{\mu,\sigma}\right)^{n}$ with those of $F_{n}$
6:     $\hat{{{\boldsymbol{\pi}}}}\leftarrow\begin{pmatrix}0.5\\ 0.5\end{pmatrix}$, $\hat{{{\boldsymbol{\mu}}}}\leftarrow\begin{pmatrix}\mu\\ 2(\hat{\mu}_{\mathcal{X}}-0.5\mu)\end{pmatrix}$, $\hat{{{\boldsymbol{\sigma}}}}\leftarrow\begin{pmatrix}\sigma\\ 0\end{pmatrix}$
7:     ${{\boldsymbol{\theta}}}\leftarrow(\hat{{{\boldsymbol{\pi}}}},\hat{{{\boldsymbol{\mu}}}},\hat{{{\boldsymbol{\sigma}}}})$
8:     $\mathbb{Q}_{N}\leftarrow\mathbb{Q}^{{{\boldsymbol{\theta}}}}$
9:     return $\mathbb{Q}_{N}$
10:end function
The GMM $\mathbb{Q}_{N}$ learned using Algorithm 3 is used for bias correction through bootstrapping, similar to the entropic risk matching approach.
4.3 Numerical illustrations
In this section, we review several methods from the literature for estimating entropic risk and show through numerical examples that they significantly underestimate it.
@@ 4.3 Numerical illustrations
We have already shown that naively fitting a GMM underestimates the true risk. In Figure 3, using the setup from example 2, we demonstrate that our methods (BS-Match and BS-EVT) help mitigate this underestimation, even with finite samples. For $N>10000$, the bias correction decays at a rate similar to the bootstrapped true bias estimate. However, EVT tends to overestimate the correction.
The next example illustrates that several approaches from the literature underestimate the true entropic risk significantly.
@@
Figure 3: Bias correction estimated from fitting a GMM by MLE, entropic risk matching (Match) and tail fitting (EVT) followed by bootstrapping.

### v3

@@ 4.1 Entropic Risk Matching
In this section, we describe Algorithm 2 which is used to learn the parameters ${{\boldsymbol{\theta}}}$ of the GMM, $\mathbb{Q}^{{{\boldsymbol{\theta}}}}$, such that the distribution of entropic risk for $n$ samples drawn from $\mathbb{Q}^{{{\boldsymbol{\theta}}}}$ matches the empirical distribution of entropic risk for $n$ samples drawn from the empirical loss scenarios $\mathcal{S}$. To construct the latter distribution, $\mathcal{S}$ is divided into $B$ bins, with each bin containing $n=N/B$ scenarios. The entropic risk is computed for each bin, forming the set $\mathcal{R}_{\mathcal{D}_{N}}$. The corresponding empirical distribution, ^$\hat{\mathbb{P}}_{\mathcal{R}_{\mathcal{D}_{N}}}$, over the set $\mathcal{R}_{\mathcal{D}_{N}}$, captures the variability of entropic risk across the $B$ bins. For the former distribution, with a fixed ${{\boldsymbol{\theta}}}$, $B^{\prime}\times n$ i.i.d. samples are drawn from $\mathbb{Q}^{{{\boldsymbol{\theta}}}}$ and divided into $B^{\prime}$ bins. The entropic risk is then computed for the scenarios in each bin, yielding the set $\mathcal{R}_{{{\boldsymbol{\theta}}}}$. The corresponding empirical distribution, ^$\hat{\mathbb{P}}_{\mathcal{R}_{{\boldsymbol{\theta}}}}$, captures the variability of entropic risk across the $B^{\prime}$ bins.
@@
Next, the algorithm compares the empirical distribution ^$\hat{\mathbb{P}}_{\mathcal{R}_{\mathcal{D}_{N}}}$ with the model-based distribution ^$\hat{\mathbb{P}}_{\mathcal{R}_{{\boldsymbol{\theta}}}}$ using the following Wasserstein distance:
[…]
@@
where $F_{\mathcal{R}_{\mathcal{D}_{N}}}^{-1}$ and $F_{\mathcal{R}_{{\boldsymbol{\theta}}}}^{-1}$ are quantile functions associated with sets $\mathcal{R}_{\mathcal{D}_{N}}$ and $\mathcal{R}_{{\boldsymbol{\theta}}}$, respectively. This distance quantifies the discrepancy between the two distributions. The algorithm iteratively adjusts the GMM parameters to minimize this distance. It uses gradient descent to update the parameters ${{\boldsymbol{\theta}}}_{t}$ at each iteration as follows:
@@
${{\boldsymbol{\theta}}}_{t+1}={{\boldsymbol{\theta}}}_{t}-\gamma\nabla_{{{\boldsymbol{\theta}}}_{t}}\mathcal{W}^{2}\left(\hat{\mathbb{P}}_{\mathcal{R}_{\mathcal{D}_{N}}},\hat{\mathbb{P}}_{\mathcal{R}_{{{\boldsymbol{\theta}}}_{t}}}\right),$
where $\gamma$ is the step size. To enable computation of the gradients of the Wasserstein distance with respect to ${{\boldsymbol{\theta}}}$, the algorithm employs differentiable sampling techniques that enable automatic differentiation through the sampling process (see Algorithm 6 in Appendix B.4). This is based on reparameterization approach (Kingma et al.,, 2015), which allows stochastic sampling operations to be expressed in a differentiable manner. The iterative process continues until the Wasserstein distance $\mathcal{W}^{2}\left(\hat{\mathbb{P}}_{\mathcal{R}_{\mathcal{D}_{N}}},\hat{\mathbb{P}}_{\mathcal{R}_{{{\boldsymbol{\theta}}}_{t}}}\right)$ falls below a predefined convergence threshold $\epsilon$, or until a maximum number of iterations $T$ is reached. Further details of the algorithm can be found in Appendix B.3.
Even though computing the Wasserstein distance between distribution of losses has a worst-case complexity $O(B^{\prime}\log(B^{\prime}))$ (Kolouri et al.,, 2019), there is a significant total cost associated with the gradient descent procedure described in Algorithm 2. In the next section, we provide a semi-analytic procedure to learn a two-component GMM that can account for the tail scenarios.
@@
1:function BS-Match($\mathcal{S}$)
2:     Divide loss scenarios in $\mathcal{S}$ into $B$ bins, each of size $n$
3:     Compute the entropic risk in $B$ bins, forming the empirical distribution ^$\hat{\mathbb{P}}_{\mathcal{R}_{\mathcal{D}_{N}}}$
4:     Fit a GMM $\mathbb{Q}^{{\boldsymbol{\theta}}}$ to $\mathcal{S}$ with Y components using EM algorithm
5:     Initialize the iteration counter $t\leftarrow 0$ and $\mathfrak{D}\leftarrow\infty$
@@
8:         Compute entropic risk in each bin, forming ^$\hat{\mathbb{P}}_{\mathcal{R}_{{{\boldsymbol{\theta}}}_{t}}}$
9:         Update GMM parameters: ${{\boldsymbol{\theta}}}_{t+1}\leftarrow{{\boldsymbol{\theta}}}_{t}-\gamma\nabla_{{{\boldsymbol{\theta}}}_{t}}\mathcal{W}^{2}\left(\hat{\mathbb{P}}_{\mathcal{R}_{\mathcal{D}_{N}}},\hat{\mathbb{P}}_{\mathcal{R}_{{{\boldsymbol{\theta}}}_{t}}}\right)$
@@
11:         Update distance: $\mathfrak{D}\leftarrow\mathcal{W}^{2}\left(\hat{\mathbb{P}}_{\mathcal{R}_{\mathcal{D}_{N}}},\hat{\mathbb{P}}_{\mathcal{R}_{{{\boldsymbol{\theta}}}_{t}}}\right)$
@@
14:     return $\mathbb{Q}^{{{\boldsymbol{\theta}}}_{t}}$
15:end function
@@ 4.2 Matching the extremes
This section focuses on learning the parameters ${{\boldsymbol{\theta}}}$ of a GMM $\mathbb{Q}^{{\boldsymbol{\theta}}}$ by giving special attention to the tails of the loss distribution, aiming to improve the accuracy of the entropic risk estimation. Although the entropic risk depends on the entire distribution, its sensitivity to extreme values makes an accurate approximation of tail behavior essential. Figure 14 in Appendix B.1 illustrates the influence function, highlighting how tail events disproportionately affect the entropic risk. Our distribution fitting approach strikes a balance between capturing the tail behavior and retaining analytical tractability.
@@
Our proposed procedure is motivated by the Fisher–Tippett–Gnedenko extreme value theorem (de Haan and Ferreira,, 2006) which states that given i.i.d. samples of $\{\zeta_{1},\zeta_{2},\cdots,\zeta_{n}\}$ with cumulative distribution function (cdf) given by $F(\cdot)$, the distribution of the (normalized) maxima $M_{n}=\mathop{\rm max}\{\zeta_{1},\zeta_{2},\cdots,\zeta_{n}\}$ converges to a non-degenerate distribution $G$:
$\lim_{n\rightarrow\infty}{\mathbb{P}}\left(\frac{M_{n}-b_{n}}{a_{n}}\leq x\right)=\lim_{n\rightarrow\infty}F(a_{n}x+b_{n})^{n}\rightarrow G(x),$
@@
where $a_{n}$ and $b_{n}$ are normalizing sequences of scale and location parameters, respectively, that ensure the limit exists and $F(\cdot)^{n}$ is the cdf of $M_{n}$. The limit distribution $G$ belongs to one of three extreme value distributions–Weibull, Fréchet or Gumbel distribution–depending on the tails of $F(\cdot)$.
Algorithm 3 constructs an equally-weighted two-component GMM to represent the loss distribution. The first component aims to estimate the distribution of maxima $M_{n}$ of the loss scenarios. To this end, first we divide the loss scenarios $\mathcal{S}$ into $B$ bins, each of size $n=N/B$. We store the maximum within each bin in set $\mathcal{M}$, where $F_{\mathcal{M}}$ denotes the cdf of scenarios in $\mathcal{M}$. This approach is typically referred to as the block maxima method (de Haan and Ferreira,, 2006). Motivated by the extreme value theory, the algorithm estimates the parameters of the first component by matching the cdf of maximum of $n$ i.i.d. samples from $\mathcal{N}(\mu^{e},\sigma^{e})$, denoted by $\Phi_{\mu^{e},\sigma^{e}}^{n}$ to $F_{\mathcal{M}}$. The calculation of the parameters $(\mu^{e},\sigma^{e})$ can be done in a semi-analytic way as discussed in Appendix B.3. For the second component of the GMM, we set the mean to $2(\mu_{\mathcal{S}}-0.5\mu^{e})$, where $\mu_{\mathcal{S}}$ is the mean of the loss scenarios, and set the standard deviation to zero. This choice ensures that the mean of overall GMM is equal to the mean of the loss scenarios $\mu_{\mathcal{S}}$.
The reasoning behind the first component of the GMM is the following. Fitting an extreme value distribution directly to the maxima $\mathcal{M}$ is computationally intensive, often inaccurate with finite samples, and typically only informative for upper tail risk, even with large datasets. Instead, fitting a normal distribution with our proposed approach is computationally efficient since it can be done semi-analytically. Furthermore, our approach can provide a reasonable approximation of the tails since both $\Phi_{\mu^{e},\sigma^{e}}^{n}$ and tails of true loss distribution converge to an extreme value distribution as $n$ goes to infinity. We emphasize that our approach is different from fitting a normal distribution directly to scenarios in $\mathcal{S}$ or $\mathcal{M}$ by MLE.
@@
1:function BS-EVT($\mathcal{S}$)
2:     Divide scenarios in $\mathcal{S}$ into $B$ bins of size $n=N/B$ each
3:     $F_{\mathcal{M}}\leftarrow$ cdf of maxima in each bin
4:     Determine $(\mu^{e},\sigma^{e})$ so that $\Phi_{\mu^{e},\sigma^{e}}^{n}$ matches $F_{\mathcal{M}}$
5:     $\hat{{{\boldsymbol{\pi}}}}\leftarrow\begin{pmatrix}0.5\\ 0.5\end{pmatrix}$, $\hat{{{\boldsymbol{\mu}}}}\leftarrow\begin{pmatrix}\mu^{e}\\ 2(\mu_{\mathcal{S}}-0.5\mu^{e})\end{pmatrix}$, $\hat{{{\boldsymbol{\sigma}}}}\leftarrow\begin{pmatrix}\sigma^{e}\\ 0\end{pmatrix}$
6:     ${{\boldsymbol{\theta}}}\leftarrow(\hat{{{\boldsymbol{\pi}}}},\hat{{{\boldsymbol{\mu}}}},\hat{{{\boldsymbol{\sigma}}}})$
7:     return $\mathbb{Q}^{{{\boldsymbol{\theta}}}}$
8:end function
@@
We end this section by revisiting Example 2, to include the estimation procedures from Sections 4.1 and 4.2. We have already shown in Figure 2 that naively fitting a GMM using MLE underestimates the true risk. In Figure 3, we observe the behavior of the bootstrap (Algorithm 1) in conjunction with the entropic risk matching (Algorithm 2) denoted by BS-Match, and the bootstrap (Algorithm 1) in conjunction with the extremes matching (Algorithm 3) denoted by BS-EVT. We can observe that the bias correction using BS-Match decays at a rate similar to the true bias. However, the bias associated with BS-MLE decays at the faster rate compared to the true bias estimate, whereas for BS-EVT, it decays slowly Thus, BS-MLE tends to underestimate the risk, while BS-EVT tends to overestimate the true risk.
@@
Figure 3: Statistics of bias correction estimated from fitting a GMM by MLE (BS-MLE), entropic risk matching (BS-Match) and tail fitting (BS-EVT) followed by bootstrapping over $100$ resampling from the underlying distribution.
4.3 Comparison with benchmarks from the literature
In this section, we review several methods from the literature for estimating entropic risk and show through a numerical example that they significantly underestimate it. We start with a brief summary of each method.
Entropic risk measure can be equivalently written as an optimized certainty equivalent risk measure (Ben-Tal and Teboulle,, 1986), i.e.,
where $h(t,\ell({{\boldsymbol{\xi}}}))=t+\frac{1}{\alpha}\exp(\alpha(\ell({{\boldsymbol{\xi}}})-t))-\frac{1}{\alpha}$.
Sample average approximation (SAA): The SAA estimator is obtained by solving SAA of problem (6):
$\rho_{\text{{SAA}}}:=\frac{1}{\alpha}\log({\mathbb{E}}_{\hat{{\mathbb{P}}}_{N}}[\exp(\alpha\ell({{\boldsymbol{\xi}}}))])=\mathop{\rm inf}_{t}{\mathbb{E}}_{\hat{{\mathbb{P}}}_{N}}[h(t,\ell({{\boldsymbol{\xi}}}))],$
where ${\mathbb{P}}$ is replaced with ^$\hat{{\mathbb{P}}}_{N}$. It is well-known that decisions based on the SAA can suffer from optimizer’s curse, leading to an optimistic bias (Smith and Winkler,, 2006). To mitigate the underestimation of the optimal value $\rho_{\mathbb{P}}(\ell({{\boldsymbol{\xi}}}))$ by the SAA estimator, leave-one out CV and optimizer’s Information Criteria (OIC) are proposed in the literature which are discussed below.
Leave-one out cross validation (LOOCV): Let ^$\hat{\mathbb{P}}_{N_{-i}}$ denote the empirical distribution without the $i$th scenario, and let $\hat{t}_{-i}$ denote the optimal solution of (6) in which ^$\hat{\mathbb{P}}_{N_{-i}}$ is used. The estimator is then defined as:
$\rho_{\mbox{\scriptsize{LOOCV}}}:=\frac{1}{N}\sum_{i=1}^{N}\left(\hat{t}_{-i}+\frac{1}{\alpha}\left(\exp(\alpha(\ell(\hat{{{\boldsymbol{\xi}}}}_{i})-\hat{t}_{-i}))-1\right)\right).$
Since $\hat{t}_{-i}$ is a feasible solution of (6), we have $\rho_{{\mathbb{P}}}(\ell({{\boldsymbol{\xi}}}))\leq{\mathbb{E}}_{{\mathbb{P}}}[h(\hat{t}_{-i},\ell({{\boldsymbol{\xi}}})]$ almost surely with respect to the randomness of $\hat{t}_{-i}$ for all $i\in[N]$. Thus,
$\rho_{{\mathbb{P}}}(\ell({{\boldsymbol{\xi}}}))$ $\,\leq\,{\mathbb{E}}\left[\frac{1}{N}\sum_{i=1}^{N}{\mathbb{E}}_{{\mathbb{P}}}(h(\hat{t}_{-i},\ell({{\boldsymbol{\xi}}})))\right]$
$=\frac{1}{N}\sum_{i=1}^{N}{\mathbb{E}}\left[{\mathbb{E}}_{{\mathbb{P}}}(h(\hat{t}_{-i},\ell({{\boldsymbol{\xi}}})))\right]$
$\,=\,\frac{1}{N}\sum_{i=1}^{N}{\mathbb{E}}\left[{\mathbb{E}}(h(\hat{t}_{-i},\ell(\hat{{{\boldsymbol{\xi}}}}_{i}))|\{\hat{{{\boldsymbol{\xi}}}}_{j}\}_{j\in[N]_{-i}})\right]$
$\,=\,\frac{1}{N}\sum_{i=1}^{N}{\mathbb{E}}[h(\hat{t}_{-i},\ell(\hat{{{\boldsymbol{\xi}}}}_{i}))]$
$\,=\,{\mathbb{E}}[\rho_{\mbox{\scriptsize{LOOCV}}}].$
where the first equality follows from linearity of expectation, the second is due to the independence of $\hat{{{\boldsymbol{\xi}}}}_{i}\sim{\mathbb{P}}$ and $\{\hat{{{\boldsymbol{\xi}}}}_{j}\}_{j\in[N]_{-i}}$ and the third follows from the law of iterated expectations. Thus, $\rho_{\mbox{\scriptsize{LOOCV}}}$ is a positively biased estimator of $\rho_{{\mathbb{P}}}(\ell({{\boldsymbol{\xi}}}))$.
Optimizer’s information criteria (OIC): To correct the first-order optimistic bias associated with the SAA, Iyengar et al., (2023) introduced the OIC estimator:
$\rho_{\mbox{\scriptsize{OIC}}}:=\rho_{\mbox{\scriptsize{SAA}}}+\frac{\text{Var}_{\hat{{\mathbb{P}}}_{N}}(\exp(\alpha\ell({{\boldsymbol{\xi}}})))}{N\alpha(\mathbb{E}_{{\hat{{\mathbb{P}}}_{N}}}[\exp(\alpha\ell({{\boldsymbol{\xi}}}))])^{2}},$
(see Appendix B.2 for detailed calculations). [^2] It is to be noted that the OIC estimator is designed to address the optimizer’s curse, while we aim to debias the empirical entropic risk estimator. The OIC estimator relies on the asymptotic normality and linearity assumption of SAA estimator. With heavy tailed losses, these assumptions on SAA estimator do not carry over to practical scenarios with limited data, hence, the underestimation issue persists. Even though the true performance of a decision obtained by solving the SAA problem overestimates the true entropic risk, the OIC estimator based on the empirical distribution underestimates the true risk.
Maximum Likelihood Estimation (MLE): The MLE estimator involves fitting a GMM $\mathbb{Q}^{{\boldsymbol{\theta}}}$ to the loss scenarios using the EM algorithm, and the estimator is given by $\rho_{\text{{MLE}}}:=(1/\alpha)\log({\mathbb{E}}_{\mathbb{Q}^{{\boldsymbol{\theta}}}}[\exp(\alpha\ell({{\boldsymbol{\xi}}}))])$.
Median-of-means (MoM): The MoM estimator is constructed by dividing the loss scenarios into $\lfloor\sqrt{N}\rfloor$ blocks, calculating the entropic risk within each block, and then taking the median of these entropic risk values (Lugosi and Mendelson,, 2019).
Bootstrap (BS): The BS estimator is calculated by repeatedly sampling $N$ observations with replacement from the empirical distribution of loss scenarios. For each bootstrap sample, the empirical entropic risk is computed. After a large number of repetitions, the average of these entropic risks provides the bootstrap estimator.